OpenStudy (anonymous):

WILL GIVE MEDAL Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3 A. x – y = y + 3 B. x – (y + 3) = 7 C. x = (7 – x) + 3 D. (y + 3) – y = 7

OpenStudy (daniellelovee):

which one do you think it is?

OpenStudy (anonymous):

C

OpenStudy (daniellelovee):

i dont think so because -y=7-x therefore positive y=-7+x or x-7

OpenStudy (anonymous):

Which one do you think it is?

OpenStudy (daniellelovee):

D

OpenStudy (daniellelovee):

thanks for the medal :)